OpenStudy (anonymous):

use the square root property to solve the quadratic equation 7x^2-35=0 a. x= +/ - sqrt of 5 b. x= sqrt of 5 c. x= + / - sqrt 35 / 7 d…

OpenStudy (anonymous):

\[7x^{2}=35\]\[x^{2}=35/7\]\[x = \pm \sqrt{ 35/7} \]

OpenStudy (anonymous):

thank you that was explained beautifuly now i understand it :)\